NEED HELP ASAP!!!!! I need help on the whole set up of negative exponents. I know it takes practice to learn it but I just don't understand how to solve them, please help.
Answer by rodriguezh2(15)

Example:
x^(-m) = (1/x^m)
So: 2^(-3) = (1/2^3) = (1/8)
Whenever you have a negative exponent you take the reciprocal of that number. So it becomes 1 divided by the number raised to the exponent
Hope this helps a little bit
If you have (a/b)^(-1) then we can say this = (b/a)
If a fraction is raised to a negative exponent we can flip the fraction
If it is raised to a power other than -1
ex: (a/b)^ (-2) then we can say this = (b^2/a^2)
(a) and (b) represent any positive number